The function A = A0e-0.00866x models the amount in pounds of a particular radioactive material stored in a concrete vault, where x is the number of years since the material was put into the vault. If 600 pounds of the material are initially put into the vault, how many pounds will be left after 120 years?
Answer by stanbon(75887)

Linear-equations/1039906 (2016-07-01 20:40:26): The function A = Ao*e-0.00866x models the amount in pounds of a particular radioactive material stored in a concrete vault, where x is the number of years since the material was put into the vault. If 600 pounds of the material are initially put into the vault, how many pounds will be left after 120 years?
---------------------------
A(x) = Ao*e(-0.00866x)
------------
A(120) = 600*e^(-0.00866*120) = 212.24 lbs
------
Cheers,
Stan H.
----------